# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### v3

@@ Abstract.
Whilst spectral Graph Neural Networks (GNNs) are theoretically well-founded in the spectral domain, their practical reliance on polynomial approximation implies a profound linkage to the spatial domain. As previous studies rarely examine spectral GNNs from the spatial perspective, their spatial-domain interpretability remains elusive, e.g., what information is essentially encoded by spectral GNNs in the spatial domain? In this paper, to answer this question, we establish a theoretical connection between spectral filtering and spatial aggregation, unveiling an intrinsic interaction that spectral filtering implicitly leads the original graph to an adapted new graph, explicitly computed for spatial aggregation. Both theoretical and empirical investigations reveal that the adapted new graph not only exhibits non-locality but also accommodates signed edge weights to reflect label consistency among nodes. These findings thus highlight the interpretable role of spectral GNNs in the spatial domain and inspire us to rethink graph spectral filters beyond the fixed-order polynomials, which neglect global information. Built upon the theoretical findings, we revisit the state-of-the-art spectral GNNs and propose a novel Spatially Adaptive Filtering (SAF) framework, which leverages the adapted new graph by spectral filtering for an auxiliary non-local aggregation. Notably, our proposed SAF comprehensively models both node similarity and dissimilarity from a global perspective, therefore alleviating persistent deficiencies of GNNs related to long-range dependencies and graph heterophily. Extensive experiments over 13 node classification benchmarks demonstrate the superiority of our proposed framework to the state-of-the-art models.
1. Introduction
@@
Graph Neural Networks (GNNs) have shown remarkable abilities to uncover the intricate dependencies within graph-structured data, and achieved tremendous success in graph machine learning (Wu et al., 2021; Chami et al., 2022; Gao et al., 2023). Spectral GNNs are a class of GNNs rooted in spectral graph theory (Chung, 1996; Shuman et al., 2013), implementing graph convolutions via spectral filters (Defferrard et al., 2016; Kipf and Welling, 2017). Whilst various spectral filtering strategies (Levie et al., 2018; Chien et al., 2021; He et al., 2021; Ju et al., 2022; He et al., 2022; Wang and Zhang, 2022; Guo et al., 2023; Tao et al., 2023; Guo and Wei, 2023b) have been proposed for spectral GNNs, their practical implementations invariably resort to approximating graph filters with fixed-order polynomials for computational efficiency (Wang and Zhang, 2022; He et al., 2022). This truncated approach essentially relies on the direct extraction of spatial features from the local regions of nodes. As such, the spatial domain of a graph, albeit loosely connected to spectral GNNs in theory, still plays a crucial role in effectively learning node representations.
@@
However, there is a notable lack of research examining spectral GNNs from the spatial perspective. Though recent studies analyze both spectral and spatial GNNs to elucidate their similarities in model formulations (Balcilar et al., 2020; Chen et al., 2021; Guo and Wei, 2023a), outcomes (Ma et al., 2021; Zhu et al., 2021), and expressiveness (Chen et al., 2020; Balcilar et al., 2021; Wang and Zhang, 2022; Sun et al., 2023), they ignore exploring the interpretability that could arise mutually from the other domain. Specifically, whilst most spectral GNNs have well explained their learned graph filters in the spectral domain (He et al., 2021; Wang and Zhang, 2022; He et al., 2022; Guo et al., 2023), understandings from the spatial viewpoint are merely limited to fusing multi-scale graph information (Liao et al., 2019); this unfortunately lacks a deeper level of interpretability in the vertex domain. Therefore, a natural question arises: what information is essentially encoded by spectral GNNs in the spatial domain?
@@
In this work, we attempt to answer this question by exploring the connection between spectral filtering and spatial aggregation. The former is the key component in spectral GNNs, while the latter is closely associated with spatial GNNs utilizing recursive neighborhood aggregation. In existing GNN frameworks, these two approaches rarely interact each other at the risk of domain information trade-offs due to uncertainty principles (Heisenberg, 1927; Folland and Sitaram, 1997; Agaskar and Lu, 2013). Recognizing the spatial significance in spectral filtering, He et al. (2021) have recently considered non-negative constraints as part of a generalized graph optimization problem. Notably, however, spatial aggregation meanwhile resembles the optimizing trajectory of the same optimization problem through iterative steps, which may be easily overlooked. Inspired by such observation, we examine, for the first time, the theoretical interaction between spectral filtering and spatial aggregation. This exploration has led us to uncover an intriguing theoretical interplay, i.e., spectral filtering implicitly modifies the original graph, transforming it into a new one that explicitly functions as a computation graph for spatial aggregation. Delving deeper, we discover that the adapted new graph enjoys some desirable properties, enabling a direct link among nodes that originally require multiple hops to do so, thereby exhibiting nice non-locality. Moreover, we find that the new graph edges allow signed weights, which turns out capable of distinguishing between label agreement and disagreement of the connected nodes.
@@
Overall, these findings underscore the interpretable role and significance of spectral GNNs in the spatial domain, inspiring us to rethink graph spectral filters beyond the fixed-order polynomials, which limit the effective propagation range of models and hinder their ability to capture long-range dependencies. Concretely, we propose a novel Spatially Adaptive Filtering (SAF) framework, for fully exploring spectral GNNs in the spatial domain. SAF leverages the adapted new graph by spectral filtering for auxiliary spatial aggregation and allows individual nodes to flexibly balance between spectral and spatial features. By performing non-local aggregation with signed edge weights, our SAF adeptly overcomes the limitations of truncated polynomials, enabling the model to capture both node similarity and dissimilarity at a global scale. As a benefit, it can mitigate persistent deficiencies of GNNs regarding long-range dependencies and graph heterophily. The contributions are summarized as follows:
@@
  • Our investigation into spectral GNNs in the spatial domain reveals that graph spectral filtering fundamentally alters the original graph, imbuing it with non-locality and signed edge weights that discern label consistency among nodes.
@@
  • Extensive experiments over 13 node classification benchmarks exhibit notable improvements of up to 15.37%, and show that SAF beats the best-performing spectral GNNs on average.
2. Related Works
2.1. Graph Neural Networks
@@
GNNs can be broadly divided into spatial-based and spectral-based models. Spatial GNNs leverage the spatial connections among nodes to perform message passing, also known as spatial aggregation (Gilmer et al., 2017; Hamilton et al., 2017). For a thorough review, we direct readers to the works (Zhou et al., 2020; Wu et al., 2021). Spectral GNNs leverage the graph’s spectral domain for convolution or, alternatively, spectral filtering (Defferrard et al., 2016; Kipf and Welling, 2017; Dong et al., 2021). Prevailing approaches focus on developing polynomial graph filters, by either learning polynomial coefficients, such as GPR-GNN (Chien et al., 2021), BernNet (He et al., 2021), ChebNetII (He et al., 2022), and JacobiConv (Wang and Zhang, 2022), or concurrently optimizing the polynomial basis for better real-world adaption, as seen in models like LON-GNN (Tao et al., 2023) and OptBasisGNN (Guo and Wei, 2023b). Diverging from this trend, ARMA (Bianchi et al., 2021) employs rational filter functions while still approximating them with polynomials. Although these methods are theoretically grounded in the spectral domain, their practical reliance on polynomial approximation hints at a profound linkage to the spatial domain. However, the spatial-domain interpretation of spectral GNNs is rarely examined. To this end, we delve into in this paper the intrinsic information spectral GNNs convey within the spatial context.
2.2. Unified Viewpoints for GNNs
@@
Several works have explored the nuances between spatial and spectral GNNs. Early studies by Balcilar et al. (2020) and Chen et al. (2021) examined their similarities in model formulations. Chen et al. (2020) proved their spatial GNN’s anti-oversmoothing ability via spectral analysis. Ma et al. (2021) and Zhu et al. (2021) utilized the graph signal denoising problem to integrate both GNN types. Balcilar et al. (2021) and Wang and Zhang (2022) further explored their expressiveness equivalence. Recently, Sun et al. (2023) have highlighted the feature space constraints of both spatial and spectral GNNs, while Guo and Wei (2023a) attempted to combine them via a residual connection module. Though these studies effectively bridge spectral and spatial GNNs, they remain focused on congruencies. Unlike them, our work represents the first endeavor to delve into the interpretability of spectral GNNs in the spatial domain, emphasizing the theoretical synergy between spectral filtering and spatial aggregation. The empirical success of our proposed method (as compared to unified GNNs in Tables 1 and 2), stemming from this in-depth analysis, further underscores our practical contributions to the literature.
2.3. Long-range Dependencies
@@
While substantial efforts have been directed towards capturing long-range dependencies in spatial GNNs (Gasteiger et al., 2019; Chen et al., 2020; Eliasof et al., 2021; Pei et al., 2020; Liu et al., 2021; Wu et al., 2022), the exploration of the same challenge in spectral GNNs remains under-studied. To fill this gap, we propose a SAF framework, which emerges as a valuable consequence of analyzing spectral GNNs in the spatial domain, enhancing their long-range dependency capture. Concurrently, Bo et al. (2023) also introduced Specformer to addresses long-range dependencies for spectral GNNs, using a Transformer based set-to-set spectral filter. However, it lacks spatial-domain interpretability and introduce more trainable parameters. In contrast, our approach creates a non-local new graph without learning additional parameters, simultaneously elucidating the interpretive implications of spectral GNNs in the spatial domain.
2.4. Graph Heterophily
Graph heterophily (Pei et al., 2020; Zhu et al., 2020b), where different labeled nodes connect, challenges GNNs operating under the homophily assumption (McPherson et al., 2001). Although many GNNs have been crafted to manage heterophilic connections (Bo et al., 2021; Wang et al., 2021; Chien et al., 2021; Yan et al., 2023; Yoo et al., 2023; Cui and Wei, 2023), our proposed SAF stands out in addressing graph heterophily. Specifically, SAF innovatively conducts an auxiliary non-local aggregation using signed edge weights, emphasizing both intra-class similarity and inter-class difference on a global scale. One should note that a recent work (Li et al., 2022) bear some resemblance to ours, introducing GloGNN and GloGNN++ to capture global homophily beyond immediate neighborhoods by learning signed edge weights. However, their approach, albeit demonstrating a grouping effect (Li et al., 2020), restricts the optimization objective into a K-hop neighborhood, focusing on similar local structural information. Conversely, our SAF framework ensures the theoretical properties of non-local learning (as proved in Section 4.2), while effectively modeling label relationships. This capability directly benefits downstream classification tasks, offering a notable superiority on real-world applications.
3. Notations and Preliminaries
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$, where the number of nodes is denoted by $N$. We define the adjacency matrix as $\mathbf{A}\in\mathbb{R}^{N\times N}$ with $A_{i,j}$ representing the edge weight between node pairs $v_{i},v_{j}\in\mathcal{V}$. The degree matrix $\mathbf{D}$ is obtained by summing the rows of $\mathbf{A}$ into a diagonal matrix. We denote the graph Laplacian matrix as $\mathbf{L}=\mathbf{D}-\mathbf{A}$, which is often normalized into $\hat{\mathbf{L}}=\mathbf{I}_{N}-\hat{\mathbf{A}}$ with $\hat{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ and $\mathbf{I}_{N}$ being an identity matrix. Let $\hat{\mathbf{L}}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$ be the eigendecomposition of $\hat{\mathbf{L}}$, where the columns of $\mathbf{U}$ refer to eigenvectors and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{N})$ consists of eigenvalues with each $\lambda_{n}\in\left[0,2\right]$. For node classification on graph $\mathcal{G}$, nodes are associated with a feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$ where $F$ refers to raw feature dimensions, and each of them is assigned a class $c_{i}$ with a one-hot vector $\mathbf{y}_{i}\in\mathbb{R}^{C}$ where $C\leq N$ is class number.
@@
Spectral Filtering. Spectral filtering is essential in spectral GNNs. It selectively shrinks or amplifies the Fourier coefficients of node features (Defferrard et al., 2016) and usually take the form as
(1) $\mathbf{Z}=g_{\psi}(\hat{\mathbf{L}})\mathbf{X}=\mathbf{U}g_{\psi}(\mathbf{\Lambda})\mathbf{U}^{T}\mathbf{X}.$
@@
Spatial Aggregation. Spatial Aggregation is a central component of spatial GNNs, facilitating the propagation of node information along graph edges and its subsequent aggregation within node neighborhood. To provide a more formal illustration of spatial aggregation, let’s consider the widely adopted GNN model, APPNP (Gasteiger et al., 2019). This model begins by applying a feature transformation, given by $\mathbf{Z}^{(0)}=f(\mathbf{X})$. The propagation then proceeds as:
(2) $\mathbf{Z}^{(k)}=(1-\eta)\mathbf{Z}^{(0)}+\eta\hat{\tilde{\mathbf{A}}}\mathbf{Z}^{(k-1)},\quad k=1,2,\cdots,K,$

### v5

@@ Abstract
Whilst spectral Graph Neural Networks (GNNs) are theoretically well-founded in the spectral domain, their practical reliance on polynomial approximation implies a profound linkage to the spatial domain. As previous studies rarely examine spectral GNNs from the spatial perspective, their spatial-domain interpretability remains elusive, e.g., what information is essentially encoded by spectral GNNs in the spatial domain? In this paper, to answer this question, we investigate the theoretical connection between spectral filtering and spatial aggregation, unveiling an intrinsic interaction that spectral filtering implicitly leads the original graph to an adapted new graph, explicitly computed for spatial aggregation. Both theoretical and empirical investigations reveal that the adapted new graph not only exhibits non-locality but also accommodates signed edge weights to reflect label consistency among nodes. These findings highlight the interpretable role of spectral GNNs in the spatial domain and inspire us to rethink graph spectral filters beyond the fixed-order polynomials, which limit the effective propagation range and hinder their ability to capture long-range dependencies. Built upon the theoretical findings, we revisit the state-of-the-art spectral GNNs and propose a novel Spatially Adaptive Filtering (SAF) framework, which leverages the adapted new graph by spectral filtering for an auxiliary non-local aggregation. Notably, our SAF comprehensively models both node similarity and dissimilarity from a global perspective, therefore alleviating persistent deficiencies of GNNs related to long-range dependencies and graph heterophily. Extensive experiments over 13 node classification benchmarks demonstrate the superiority of our proposed framework to the state-of-the-art methods.
Index Terms:
Graph Neural Networks, Spectral Filtering, Long-range Dependency, Graph Heterophily
I Introduction
@@
Graph Neural Networks (GNNs) have shown remarkable abilities to uncover the intricate dependencies within graph-structured data, and achieved tremendous success in graph machine learning [1, 2, 3, 4, 5, 6]. Spectral GNNs are a class of GNNs rooted in spectral graph theory [7, 8], implementing graph convolutions via spectral filters [9, 10]. Whilst various spectral filtering strategies [11, 12, 13, 14, 15, 16, 17, 18, 19] have been proposed for spectral GNNs, their practical implementations invariably resort to approximating graph filters with fixed-order polynomials for computational efficiency [16, 15]. This truncated approach essentially relies on the direct extraction of spatial features from the local regions of nodes. As such, the spatial domain of a graph, albeit loosely connected to spectral GNNs in theory, still plays a crucial role in effectively learning node representations.
@@
However, there is a notable lack of research examining spectral GNNs from the spatial perspective. Though recent studies analyze both spectral and spatial GNNs to elucidate their similarities in model formulations [20, 21, 22], outcomes [23, 24], and expressiveness [25, 26, 16, 27], they neglect the interpretability that could arise mutually from the other domain. Specifically, while most spectral GNNs have well explained their learned filters in the spectral domain [13, 16, 15, 17], understandings from the spatial viewpoint are merely limited to fusing multi-scale graph information [28]; this unfortunately lacks a deeper level of interpretability in the vertex domain. Therefore, a natural question arises: what information is essentially encoded by spectral GNNs in the spatial domain?
@@
In this work, we attempt to answer this question by exploring the connection between spectral filtering and spatial aggregation. The former is the key component in spectral GNNs, while the latter is closely associated with spatial GNNs utilizing recursive neighborhood aggregation. In existing GNN frameworks, these two approaches rarely interact each other at the risk of domain information trade-offs due to uncertainty principles [29, 30, 31]. Recognizing the spatial significance in spectral filtering, [13] have recently considered non-negative constraints as part of a generalized graph optimization problem. Notably, however, spatial aggregation meanwhile resembles the optimizing trajectory of the same optimization problem through iterative steps, which may be easily overlooked. Inspired by such observation, we examine, for the first time, the theoretical interaction between spectral filtering and spatial aggregation. This exploration has led us to uncover an intriguing theoretical interplay, i.e., spectral filtering implicitly modifies the original graph, transforming it into a new one that explicitly functions as a computation graph for spatial aggregation. Delving deeper, we discover that the adapted new graph enjoys some desirable properties, enabling a direct link among nodes that originally require multiple hops to do so, thereby exhibiting nice non-locality. Moreover, we find that the new graph edges allow signed weights, which turns out capable of distinguishing between label agreement and disagreement of the connected nodes.
@@
Overall, these findings underscore the interpretable role and significance of spectral GNNs in the spatial domain, inspiring us to rethink graph spectral filters beyond fixed-order polynomials, which, albeit efficient, limit models’ effective propagation range and hinder their ability to capture long-range dependencies. Concretely, we propose a novel Spatially Adaptive Filtering (SAF) framework to fully explore spectral GNNs in the spatial domain. SAF leverages the adapted new graph by spectral filtering for auxiliary spatial aggregation and allows individual nodes to flexibly balance between spectral and spatial features. By performing non-local aggregation with signed edge weights, SAF adeptly overcomes the limitations of truncated polynomials, enabling the model to capture both node similarity and dissimilarity at a global scale. As a benefit, it can mitigate persistent deficiencies of GNNs regarding long-range dependencies and graph heterophily. The contributions are summarized as follows:
@@
  • Our investigation into spectral GNNs in the spatial domain reveals that they fundamentally alters the original graph, introducing non-locality and signed edge weights to discern node label consistency.
@@
  • Extensive experiments in node classification exhibit notable improvements of up to 15.37%, and show that SAF beats the best-performing spectral GNNs on average.
II Motivation and Related Works
This section outlines the motivation behind our research, derived from thoughtful consideration of the existing related works, and elucidates how our work diverges from and contributes to the current body of research.
II-A Graph Neural Networks
@@
GNNs can be broadly divided into spatial-based and spectral-based methods. Spatial GNNs leverage the spatial connections among nodes to perform message passing, also known as spatial aggregation [32, 33] (readers are directed to works [34, 2] for a thorough review). Spectral GNNs leverage the graph’s spectral domain for convolution or, alternatively, spectral filtering [10, 35]. Prevailing approaches focus on developing polynomial graph filters, by either learning polynomial coefficients, such as GPR-GNN [12], BernNet [13], ChebNetII [15], and JacobiConv [16], or optimizing the polynomial basis for better adaption, as seen in models like LON-GNN [18] and OptBasisGNN [19]. Diverging from this trend, ARMA [36] employs rational filter functions while still approximating them with polynomials. Although these methods are theoretically grounded in the spectral domain, their practical reliance on polynomial approximation hints at a profound linkage to the spatial domain. However, the spatial-domain interpretation of spectral GNNs is rarely examined. To this end, we delve into in this paper the intrinsic information spectral GNNs convey within the spatial context.
II-B Unified Viewpoints for GNNs.
@@
Several works have explored the nuances between spatial and spectral GNNs. Early studies by [20] and [21] examined their similarities in model formulations. [25] proved their spatial GNN’s anti-oversmoothing ability via spectral analysis. [23] and [24] utilized the graph signal denoising problem to integrate both GNN types, and their expressiveness equivalence is further explored in works [26, 16]. Recently, [27] have highlighted the feature space constraints of both spatial and spectral GNNs, while [22] tend to emphasize their relationship via residual connection. Though these studies effectively bridge spectral and spatial GNNs, they remain focused on congruencies. Unlike them, our work represents the first endeavor to delve into the interpretability of spectral GNNs in the spatial domain, emphasizing the theoretical synergy between spectral filtering and spatial aggregation. The empirical success of our method (as compared to unified GNNs in Tables II and III), stem from this in-depth analysis, further underscoring our practical contributions to the literature.
II-C Long-range Dependencies
@@
While substantial efforts have been directed towards capturing long-range dependencies in spatial GNNs [37, 25, 38, 39, 40, 41, 42], the exploration of the same challenge in spectral GNNs remains under-studied. Specifically, most spectral GNNs approximate graph filters with fixed-order polynomials, which, albeit efficient, limits the effective propagation range and hinder their ability to capture the long-range dependencies. To fill this gap, we propose SAF framework that emerges as a valuable consequence of analyzing spectral GNNs in the spatial domain. Aligned with our objective, Specformer [43] is introduced to addresses long-range dependencies for spectral GNNs, using a Transformer based set-to-set spectral filter. However, it lacks spatial-domain interpretability and introduce more trainable parameters. In contrast, our approach creates a non-local new graph without learning additional parameters, meanwhile elucidating the spatial implications of spectral GNNs. Similarly, a recent approach, FLODE [44], also produces matrices capable of capturing long-range dependencies by utilizing the fractional graph Laplacian [45]. However, this work mainly generalizes existing concepts from undirected to directed graphs to mitigate graph oversmoothing. Our study, conversely, focuses on exploring the fundamental issues surrounding undirected graphs to delve deeper into the intrinsic significance of spectral GNNs.
II-D Graph Heterophily
Graph heterophily [39, 46], where different labeled nodes connect, challenges GNNs operating under the homophily assumption [47]. Although many GNNs have been crafted to manage heterophilic connections [48, 49, 12, 50, 51, 52, 53, 54, 55], our method offers a distinct solution. Specifically, SAF innovatively conducts an auxiliary non-local aggregation using signed edge weights, emphasizing both intra-class similarity and inter-class difference on a global scale. One should note that there are two recent works [56, 57] also employ signed edge weights, introducing GloGNN and LRGNN, respectively. GloGNN aims to capture global homophily but is limited to K-hop neighborhood information, while LRGNN extends this by using low-rank properties to approximate the true global relationships between node labels. Despite their success, both methods focus primarily on fitting node label relationships without fully exploring the fundamental GNN mechanisms, leading to a heavy reliance on label supervision (Section VI-I). Differently, our SAF leverages cross-domain insights into GNNs, ensuring the theoretical soundness of non-local learning, as proven in Section IV-B. This allows SAF to consistently perform well under both sparse and dense supervision.
III Notations and Preliminaries
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with node set $\mathcal{V}$ and edge set $\mathcal{E}$, where the number of nodes is denoted by $N$. The adjacency matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$ defines the edge weights $A_{i,j}$ between nodes $v_{i}$ and $v_{j}$, and the degree matrix $\mathbf{D}$ can be obtained by summing the rows of $\mathbf{A}$ into a diagonal matrix. We denote the graph Laplacian matrix as $\mathbf{L}$, which is often normalized into $\hat{\mathbf{L}}=\mathbf{I}-\hat{\mathbf{A}}$ with $\hat{\mathbf{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}$ and an identity matrix $\mathbf{I}$. Let $\hat{\mathbf{L}}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{T}$ be the Laplacian eigendecomposition, where $\mathbf{U}$ is eigenvector matrix and $\mathbf{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{N})$ consists of eigenvalues within $\left[0,2\right]$. For node classification, we usually have a feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$ with $F$ being raw feature dimensions, and each node is assigned a one-hot label vector $\mathbf{y}_{i}\in\mathbb{R}^{C}$ where $C\leq N$ is class number.
III-A Spectral Filtering
@@
Spectral filtering is essential in spectral GNNs. It selectively shrinks or amplifies the Fourier coefficients of node features [9] for learning tasks and usually take the form as
$\mathbf{Z}=g_{\psi}(\hat{\mathbf{L}})\mathbf{X}=\mathbf{U}g_{\psi}(\mathbf{\Lambda})\mathbf{U}^{T}\mathbf{X}.$ (1)
@@
III-B Spatial Aggregation
Spatial aggregation is a central component of spatial GNNs, facilitating the propagation of node information along edges and its subsequent aggregation within node neighborhood. To provide a better illustration, let’s consider a popular spatial GNN, APPNP [37]. This model begins with a feature transformation – $\mathbf{Z}^{(0)}=f(\mathbf{X})$, and then perform the spatial aggregation as:
$\mathbf{Z}^{(k)}=(1-\eta)\mathbf{Z}^{(0)}+\eta\hat{\tilde{\mathbf{A}}}\mathbf{Z}^{(k-1)},\quad k=1,2,\cdots,K,$ (2)